A cylinder with the height of 15 has a total surface solution area that is twice the area of a circle whose radius is 6 more than the radius of the cylinder. What is the radius of the cylinder? May you help me with this> I am very lost.
Answer by scott8148(6628)

the surface area of the cylinder is the lateral area (2*pi*r*15) plus the area of the 2 ends (2*pi*r^2)

30*pi*r+2*pi*r^2=2*pi*(r+6)^2 ... dividing by 2*pi gives 15r+r^2=(r+6)^2

15r+r^2=r^2+12r+36 ... 15r=12r+36 ... 3r=36 ... r=12